Please, Complete questions 1, 3 and 5.

  1. Table 5.1 indicates that the average annual rate of return on common stocks over many years has exceeded the return on government bonds in the United States. Why do we observe this pattern?

2. Suppose the realized rate of return on government bonds exceeded the return on common stocks one year. How would you interpret this result?

3. What is most important to investors: the number of a company’s shares they own, the price of the company’s stock, or the value of their shareholdings in the company? Why?

4. Two 20-year bonds are identical in all respects except that one allows the issuer to call the bond in return for $1,000 cash at any time after five years, while the other contains no call provisions. Will the yield to maturity on the two bonds differ? If so, which will be higher? Why?

5. The return an investor earns on a bond over a period of time is known as the holding period return, defined as interest income plus or minus the change in the bond’s price, all divided by the beginning bond price. What is the holding period return on a bond with a par value of $1,000 and a coupon rate of 6 percent if its price at the beginning of the year was $1,050 and its price at the end was $940? Assume interest is paid annually. Can you give two reasons the price of the bond might have fallen over the year?

Table 5.1 Rate of Return on Selected U.S. Securities, 1928-2016

Security

Return

Common Stocks

11.4%

Long-term corporate bonds

5.7%

Long-term government bonds

5.2%

Short-term government bills

3.5%

Consumer price index

3.1%

Real return=5.2-3.1=2.1%(Exact real return=(1.052)/(1.031)-1=2.0%)

Default premium=5.7%-5.2%=0.5%
Answer #1

Why do we observe this pattern?

Different securities have different level of risks associated with them. The returns offered by them are commensurate with the risk profile of the instruments. Common stock has highest risk and hence they offer the highest return. Short term government bills are nearly risk free and hence they have the lowest return.

2. Suppose the realized rate of return on government bonds exceeded the return on common stocks one year. How would you interpret this result?

  • The result will be treated as exceptional and outlier.
  • The result will be surprising and astonishing as this is against the rules of corporate finance
  • This will be treated as a short term phenomenon with a hope that such an anomaly will be taken care of in the long run

3. What is most important to investors: the number of a company’s shares they own, the price of the company’s stock, or the value of their shareholdings in the company? Why?

The value of the shareholding is important. A shareholder should be concerned with the overall value enhancement. A stock split may reduce the price to half and double the number of shares, but the total value of the holding will still remain the same, thus leading to no accretion. Hence, the value of their shareholdings in the company is important for the shareholders.

4. Two 20-year bonds are identical in all respects except that one allows the issuer to call the bond in return for $1,000 cash at any time after five years, while the other contains no call provisions. Will the yield to maturity on the two bonds differ? If so, which will be higher? Why?

Yes, the yield to maturity on the two bonds differ. The bond with the call option will have a higher yield than the straight bond. The call option brings an additional value to the issuer, as the issuer has the flexibility to call the bond. Hence, the yield on the callable bonds are higher.

5. The return an investor earns on a bond over a period of time is known as the holding period return, defined as interest income plus or minus the change in the bond’s price, all divided by the beginning bond price. What is the holding period return on a bond with a par value of $1,000 and a coupon rate of 6 percent if its price at the beginning of the year was $1,050 and its price at the end was $940? Assume interest is paid annually. Can you give two reasons the price of the bond might have fallen over the year?

Holding period return = (P1 + Interest - P0) / P0 = (940 + 1,000 x 6% - 1,050) / 1,050 = -4.76%

Reasons for price fall:

  • Increase in interest rate: An increase in interest rate leads to a decline in the value of the bond
  • Fall in credit quality of the issuer: If the credit quality declines, the bonds issued by the issuer becomes cheaper